2x+y-5z+11=0
3x+y+2z-13=0
x+2y-5z+11=0

Answer by Alan3354(69443)

2x+y-5z+11=0
3x+y+2z-13=0
x+2y-5z+11=0
-------------
--> (0,0,0)